*œBO8aKœ(B

œBJBNsO@M7?œB0l4DH7F œ(BICOT (œB?7@BeœBO@œBH7F:NMQ5l?œ(B, œBHFMQNJBNsœBG3 + H/5l?JBNsO@M7?œBœB”kœ(B.

œBK%F œBACCN¡1H4pK∗J

KL1 œBN œB4pK∗J œB7+jJV7N@)8fœ(B œBœB KL1 œBNJ8K!5,B’r > $7/2r@b9k;vO7J

, œ(B, œ

BœBB?/MQ

F@bL@7F

/?aœ(B, œB < +A3Kœ(B KL1 œBN4pK5G0OM2rG−kO:G”œ(B œBkœ(B.

œBBhœ(B2œBIt=i5iJTOKJE∗JœBL∗H7œ(B, œB0J2¡N; v9‘r3XV œ(B. œBœB@aNA ∗ BrJ < 0N?’!9J œBAH9 _{=}R8lrMQ

?3F¡o œB œBBhœ(B3œBItCf5iJTO8zN(NNI

œ

B=,F@rL∗H7œ(B, œB0J2¡N; v9‘r3XV œ(B. œBAq1‘5!G =H=N;H

Jœ(Bœ

B7W;;N = ∗N;NH=Djœ(B œBMœB œBMW 5a6nF 07?œB œBC10l; 2 > Hœ(B œBJ∗œ(B, œBœB+;khJCF

kœ(B.œ

B?œ(B, œBF = P NOJINo:+JJQ99Gœ(B, œBœBO5¡;wœBK œBF LgJT œ(B